OpenStudy (anonymous):

I NEED HELP PLEASE! When one half of a number is added to one third of a number, the answer is 30. Find the number.

OpenStudy (anonymous):

36

OpenStudy (anonymous):

ok, so lets call this mystery number 'x' we know that (1/2)x + (1/3)x = 30

OpenStudy (anonymous):

I know the answer! I need to EXPLAIN how i got the answer.

OpenStudy (anonymous):

ok, so from there, we can add the fractions as long as we have a common denominator right? so lets make our common denominator 6 since that is 2 times 3

OpenStudy (anonymous):

okay then

OpenStudy (anonymous):

that means that we need to multiply the (1/2)x by (3/3) and the (1/3)x by (2/2) this will give us 3x/6 + 2x/6 = 30

OpenStudy (anonymous):

Can you write all in once? that's gonna make me easy to understand :(

OpenStudy (anonymous):

now we can add them together and we get 5x/6 = 30

OpenStudy (anonymous):

yup, i got you, here it is: (1/2)x + (1/3)x = 30 (3/3)(1/2)x + (2/2)(1/3)x = 30 3x/6 + 2x/6 = 30 5x/6 = 30 5x = 30(6) 5x = 180 x= 180/5 x = 36

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

AWWHH! Thanks a LOT :)

OpenStudy (anonymous):

no problem! hope that is clear enough!

OpenStudy (anonymous):

Yes. It is. :)